skysky2:

Subtract the following polynomials, then place the answer in the proper location on the grid. Write the answer in descending powers of x. 8.3x^2+0.9xy+3.6 7.4x^2-1.7xy+6.4

darkknight:

to subtract polynomials, you have to combine like terms. For example, if we have the functions a(x) and s(x), and we want to subtract these 2 functions we will do a(x)-s(x). Suppose a(x) = x^2+2x and s(x) is 2x^2+4x Than a(x)-s(x) = ((x^2+2x) - (2x^2+4x)) which equals -x^2-2x. So use the same process to find the answer to your problem and lmk if you need help :)